Find the sum of first 35 terms of the series whose pth term is p/7+2.

Answer by stanbon(75887)

: Find the sum of first 35 terms of the series whose pth term is p/7+2.
---------
a(1) = (1/7)+2 = 15/7
a(2) = (2/7)+2 = 16/7
a(3) = (3/7)+2 = 17/7
-----
a(35) = (35/7)+2 = 7
---------
Arithmetic Sequence with a(1) = 15/7 and d = 1/7
-------------------------
Sum:: S(35) = (35/2)((15/7)+7) = 160
---------------
Cheers,
Stan H.
---------------